Reports 1-1 of 1 Clear search Modify search
AdV-SVS (Pre-commissioning)
garaventa, mehmet, sequino, sorrentino, vahlvruch - 12:22 Monday 31 May 2021 (51978) Print this report
Squeezing measurement with delay line on EQB1 - extra losses and issue with coherent control

On Friday morning the AEI source was switched to squeezing mode. We engaged the automatic alignment on the homodyne detector to optimize the relative alignment of CC and LO beams, then we optimized the centering of HD photodiodes by translating the phododiodes to minimize the signal at M4 angular dither lines; we then engaged the coherent control with the usual procedure. We did not manage to increase the gain beyond 4000 with filter 1 without inducing an high frequency oscillation, and the UGF was limited to about 100 Hz in such conditions. We also maximized the LO power by acting on the waveplate before the PBS on LO path.

With the CC engaged we scanned the 4 MHz phase; we modified the proportional gain along the phase scan according to the change of error signal slope. Looking at the spectrum of HD audio channel, we found a maximum squeezing at 2.3 rad CC phase, and maximum anti-squeezing around 0.9 rad CC phase. We then tuned the LO telesope looking at the 4 MHz beat magnitude to slightly improve the mode matching. The measured squeezing level was about 1.7 dB and anti-squeezing was about 3.2 dB, see attached plot. We will need to search for extra losses on the SQZ path and to understand how the TF for CC loop has changed.

Images attached to this report
Comments to this report:
garaventa, sorrentino - 0:07 Tuesday 01 June 2021 (51993) Print this report

Today we did a quick check of the optical losses on the IR path by measuring the power of SC beam with a powermeter. We found that the tranmission of the Faraday Isolator on the main SQZ path is around 90%. As the polarisation was recently tuned, and the measured throughput was around 99.5% after installation on SQB1, we suspect that some recent activities on the bench might have produced some beam clipping. We'll repeat the test and possibly adjust the FI alignment with the BAB.

garaventa, sequino, delaurentis - 18:56 Wednesday 09 June 2021 (52096) Print this report
This afternoon we investigated the losses due to FI on EQB1 bench. Using BAB (SC shuttered) we measured the power before and after FI:
before 1.42mW
after 1.40mW

so we have losses due to FI of about 1.4%.

Then, we have verified that the SC is not well aligned in the FI as we measure about 10% of power losses (before the FI and in transmission to it towards the OPO).

With only BAB, we also checked the power after the FI along the path to the HD, not measuring losses.

garaventa - 17:15 Thursday 17 June 2021 (52181) Print this report
I estimated the phase noise related to the squeezing measurement on EQB1 on May 28th.
Figure 1 shows the calibration signal (gps = 1306251352).
Figure 2 shows the CC residual phase noise (during sqz measurement, gps = 1306247219). So, the phase noise is about 156 mrad.
In Figure 3 I compared the phase noise in the sqz measurement on May 28th and on February (light blue, about 187mrad).
Images attached to this comment
Search Help